
At 22:00 image 280 finished normally. Then everything just stopped but with no indication anything had happened. After a few minutes Manual pushed sop and go in the Exposure queue wiht no apparent effect. Then he reconfigured SISPI on the console. 281 was aborted and 282 came into the queue and was executed.
| Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
|---|---|---|---|---|---|---|
| 2019-03-05 | 9.5 | 0.0 | 0.0 | 0.0 | 0.0 | 9.5 |
Links
[1] mailto:dschlegel@lbl.gov
[2] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/8868/delete?destination=print/node/15385
[3] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/8868/edit?destination=print/node/15385
[4] mailto:ssheppard@carnegiescience.edu
[5] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/8869/delete?destination=print/node/15385
[6] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/8869/edit?destination=print/node/15385
[7] http://www.ctio.noao.edu/noao/field-collection/field-nr-prog/add/node/15385?destination=print/node/15385
None.